OpenStudy (anonymous):

Two integers from 1 through 40 are chosen by a random number generator. What is the probability that a) the numbers are both even, b) one number is even and one number is odd, c) both numbers are less than 30, d) the same number is chosen twice? I got a) 0.244 b) 0.53 c)0.526 (I did 29/40 x 29/40) d) ?

OpenStudy (anonymous):

I think d is 1/40*1/40

OpenStudy (anonymous):

thats waht i thought, but i looked online and i got different answers :(

OpenStudy (anonymous):

i aso need to know if i did c correctly

OpenStudy (anonymous):

Nvm I solved it, the correct answer is 40/1600 or 1/40.

OpenStudy (anonymous):

Just need someone to check my answers

OpenStudy (welshfella):

each time the generator chooses a number there are a total of 40 numbers so (a ) would be 1/2 * 1/2 = 0.25

OpenStudy (anonymous):

I dont understand how you got that, I thought the answer for d was 0.0025

OpenStudy (anonymous):

a 25% chance in the senario doesnt even make sense to me

OpenStudy (welshfella):

I'm talking about a)

OpenStudy (anonymous):

I know I got a right it is (20C2/40C2) = 190/780 = 0.2435...

OpenStudy (welshfella):

oh If the book says its right... I dont understand that. I figured that because it was a random number generator it would have 40 numbers to pick from on each draw. 0.244 is correct if , on the second draw, there is only 39 numbers to pick from.. So i guess that must be the case.

OpenStudy (anonymous):

can you check c and d for me

OpenStudy (welshfella):

sorry I can't get my head around this one. Take d :- if the first number is drawn is not included in the second draw the its impossible. If it is then the answer is 1/40 * 1/40 = 1/1600

OpenStudy (anonymous):

thats what i thought also, but looked at a similar problem online withe case being 60 instead of 40 and i did it they way they showed

OpenStudy (anonymous):

lol this is confusing me also. I think its supposed to be done one way and i see people doing it another way with a different supposedly correct answer

OpenStudy (anonymous):

I think its becasue the numbers arnt being taken one at a time. The generator is generating two numbers each 1 through 40. It's not the case of draw one, put it back,and draw another. Or draw one and then draw another.

OpenStudy (welshfella):

but how can the generator take out 2 of the same number?

OpenStudy (anonymous):

Its seperate. It generates two numbers at the same time each having the set integers of 1 through 40.

OpenStudy (welshfella):

i would try 1/1600 for d

OpenStudy (welshfella):

have you got the book answer for c)?